# Getting Python Types Right with RightTyper

### v2

@@ Abstract.
This paper presents RightTyper, a novel hybrid approach for Python that produces accurate and precise type annotations grounded in actual program behavior. RightTyper grounds inference in types observed during actual program execution and combines these observations with static analysis and name resolution to produce substantially higher-quality type annotations than prior approaches. Through principled, statistically guided adaptive sampling, RightTyper balances runtime overhead with the need to observe sufficient execution behavior to infer high-quality type annotations. We evaluate RightTyper against static, dynamic, and AI-based systems on both synthetic benchmarks and real-world code, and find that it consistently achieves higher semantic similarity to ground-truth and developer-written annotations, respectively, while incurring only approximately 25% runtime overhead.
@@ 1.1. Contributions
This paper proposes RightTyper, a novel hybrid type inference approach for Python. RightTyper sidesteps pitfalls of static and AI-based methods by sampling types observed during actual program execution. It employs adaptive sampling strategies to minimize runtime overhead while maintaining inference accuracy. RightTyper monitors function calls using a Poisson process, in which observations occur only during brief, randomly timed capture windows, allowing most instrumentation to be disabled outside these windows. For container types, RightTyper fully scans small containers, but applies Good–Turing estimation as a stopping criterion for larger ones, an estimator developed by Turing to estimate unseen events (good-turing; vanderWalt2021). Additionally, RightTyper tracks previously seen containers to avoid re-typing them, while continuing to sample to detect changes. Together, these strategies achieve low overhead (overall $\approx$25%) while using statistically principled sampling to obtain representative coverage.
@@
  • It introduces RightTyper, a novel dynamic–static hybrid type inference approach for Python (§2);
  • It presents a prototype implementation of RightTyper, released as open-source software on GitHub (§3);
  • It presents an enhanced implementation of TypeSim—a semantic type similarity metric proposed by prior work (typybench)—that addresses various issues in the original implementation and is also released on GitHub (§4.1);
  • It empirically compares the types inferred by RightTyper to those from prior work, demonstrating a marked improvement in the quality of annotations (§4.2, §4.3);
  • It evaluates RightTyper’s runtime overhead against other inference tools, showing that it is significantly more efficient, and fast enough for practical use (§4.4).
@@ 2.1. Instrumentation and Event Sampling
Most of RightTyper’s instrumentation utilizes sys.monitoring, a low-overhead, fine-grained mechanism introduced in Python 3.12 for tracking various events during execution (pep669). This feature allows selective activation and deactivation of monitoring for specific portions of user code. To capture events not supported by sys.monitoring, RightTyper additionally instruments the code (§3).
@@ 2.3. Typing Generics
For large containers, after collecting a minimum number of samples (default 32), RightTyper computes the singleton ratio $\hat{r}$—the proportion of observed types that appear exactly once:
@@
where $C$ is a counter mapping each observed type $t$ to its count $C(t)$, $n$ is the number of samples, and $\mathbf{1}[\cdot]$ is the indicator function. When this ratio falls below a threshold $\tau$ (default $\tau=0.05$) for all parameters (e.g., dictionary key and value types are counted separately), sampling terminates. A hard cap (default 64 samples) further prevents pathological cases from dominating execution time. RightTyper aggregates all types observed across inspections into a cumulative history, from which it generates annotations.
@@ 2.6. Simplifying Types
Despite the simplifications above, annotations can still be complex. RightTyper offers options to trade precision for readability, such as collapsing repeated argument types (e.g., tuple[int, int, int] to tuple[int, ...]) and limiting nesting depth.
@@ 3. Implementation
RightTyper uses sys.monitoring events (PY_START, PY_RETURN, PY_YIELD, and PY_UNWIND) to capture function arguments, return values, yielded values, and exception-driven unwinding. Yielding temporarily suspends execution to produce a value before later resuming. Generator functions and coroutines (including async functions) both yield values, but coroutines may also receive values during execution; because sys.monitoring does not expose these received values, RightTyper supplements monitoring with AST-based instrumentation, intercepting code loading to inject the necessary hooks before execution.
@@ 4. Evaluation
Do type annotations generated by RightTyper improve upon those from prior approaches? (§4.2, §4.3)
@@
How does RightTyper’s runtime overhead compare to that of other dynamic approaches? (§4.4)
@@ 4.1. Metric
The original implementation builds upon mypy and relies on the types it reports to compute similarity; this includes types inferred by mypy even in the absence of annotations, as well as types mapped to the Any wildcard when mypy cannot determine them (e.g., for dependencies lacking annotations) or abandons analysis due to earlier errors—which can either inflate or deflate TypeSim scores. We therefore reimplement TypeSim without relying on mypy. In addition, we correct an issue in the original implementation that compared only attributes declared directly on each type, yielding a TypeSim score of 0 between a type and its parent class.
We further extend TypeSim to support additional typing constructs—such as Callable, Literal, Type, and TypeAlias—which the original implementation ignored, and to better handle matching Any and object in generic parameters. Developers commonly use Any or object to avoid overly complex annotations (e.g., dict[str, Any] instead of dict[str, int|str|list[str]]).
@@
Figure 10(a) presents the results in terms of TypeSim scores. With an overall 99.8% TypeSim, RightTyper achieves a near-perfect score, with only two cases failing to yield an exact match. In one case, the benchmark requests the type of an abstract method that does not execute. In the other case, RightTyper correctly types a zip iterator using the Iterator protocol—necessary to specify element types, since zip itself does not support them—but receives a lower score. This lower score occurs because TypeSim penalizes protocol-based annotations: protocols expose only the minimal set of required operations and therefore appear less similar to richer concrete types under the metric. GPT-4o, at 96.7%, also performs well, only incorrectly inferring “Unknown” or Any in a few cases. It is followed by pytype at 60.9%; it suffers from low coverage, consistent with prior work showing that its conservative static analysis leads to limited coverage (10.1145/3652153). MonkeyType and QuAC achieve 32.9% and 23.1%, respectively, in large part because they lack support for inferring variable types (Figure 11). QuAC operates by collecting sets of required attributes and using BM25—a term-frequency ranking function from information retrieval—to match these sets to candidate types (bm25). It is therefore plausible that the benchmark’s short programs—designed to facilitate manual ground-truth verification—provide limited context for this matching process, reducing QuAC’s effectiveness (10.1145/3689783). Similarly, Type4Py’s relatively low performance (24.7%) may be due to the benchmark programs likely representing out-of-distribution inputs for its model.
@@
(b) Real-world code (6,168 annotations)
Figure 10. [RQ1] RightTyper achieves the highest type inference accuracy: (a) On TypeEvalPy, RightTyper achieves 99.8% TypeSim, outperforming GPT-4o (96.7%). (b) On real-world code, RightTyper achieves 76.8% TypeSim, outperforming GPT-4o (60.6%) and MonkeyType (52.8%).
@@
Although the programs in TypeEvalPy provide broad coverage of Python language features, they are significantly simpler than, and differ markedly from, commonly deployed Python code. We therefore next evaluate RightTyper on the real-world Python packages pylint and black, which together contain over 6,000 annotations. We select these packages based on several criteria. First, both are among the most extensively annotated open-source Python projects, with annotations developed and refined by expert maintainers over many years of use with type checkers—making them ideal ground-truth sources. Second, they are complex, production-quality codebases that exercise diverse Python features including inheritance hierarchies, generic types, protocols, decorators, context managers, and AST manipulation. Third, they are widely used (pepy.tech reports billions of downloads) and well-maintained, reducing the risk of ground-truth errors. Finally, as command-line tools, we can drive them end to end with varied inputs, achieving broader coverage than test suites alone. As ground truth, we use the projects’ original type annotations; that is, we evaluate the semantic similarity of inferred annotations to developer-written ones using TypeSim.
@@
Prior to evaluating each system, we remove all type annotations from the source code, except for special constructs such as dataclasses and NamedTuple where removing annotations would alter program semantics; in these cases, we replace the annotations with a type alias placeholder, allowing us to distinguish them from tool-inferred types during evaluation. We also remove any docstrings, as these can influence AI-based type inference; this ensures a level comparison with non-AI methods that cannot leverage natural language hints.
@@
When evaluating RightTyper and MonkeyType, we execute each target package multiple times with varying options and inputs—keeping them identical across the two systems—and also run their test suites to improve coverage, thereby collecting typing information across a range of execution scenarios. As before (§4.2), we run MonkeyType without sampling, allowing it to record every execution and achieve its best possible results. Rather than having MonkeyType annotate the source code directly, we configure it to generate stub files, as in-source annotation causes MonkeyType to emit import statements that introduce circular dependencies. Applying MonkeyType and QuAC to these sources uncovered several bugs, including an assertion failure in QuAC and syntax errors in MonkeyType’s output; we corrected these issues. To include LLM-based results, we reimplement querying GPT-4o using TypyBench’s “single-file context” prompt, as its replication package does not provide an implementation.
@@
Figure 10(b) presents the results. With an overall TypeSim score of 76.8%, RightTyper achieves the highest similarity to developer-annotated types. It is followed by GPT-4o and MonkeyType, which at 60.6% and 52.8% also show good performance, and at a considerable distance by Type4Py at 30.8%, QuAC at 26.7%, and pytype at 23.2%.
@@
Figure 11. [RQ1] Detailed results: (a) TypeSim scores by annotation category show RightTyper excels on both functions and variables, while QuAC and MonkeyType cannot infer variable types. (b) On real-world code, RightTyper achieves the highest scores across all metrics: TypeSim (76.8%), Exact Match (67.5%), and Coverage (89.8%).
@@
RightTyper exhibits some similar limitations, including occasional overly concrete typing and missed coverage, the latter being expected given that it executes the same scenarios as MonkeyType. However, RightTyper achieves higher scores overall due to more complete language support—such as inferring variable types, handling keyword and variadic arguments, and inferring Callable argument types, all of which MonkeyType lacks—as well as its ability to better simplify inferred types, yielding cleaner annotations that more closely match the originals (Figure 11). The lower scores on variable annotations stem from variables captured only at function exit (missing intermediate types), unexecuted branches, and developers annotating with broader types than observed at runtime. Finally, in some cases RightTyper correctly emits annotations with type parameters (cf. §2.4), but TypeSim yields a zero score because it does not capture these forms of semantic equivalence.
Given RightTyper’s strong results on both TypeEvalPy and real-world code, we conclude that it substantially improves annotation quality relative to prior approaches.
[RQ1] Summary: With near-perfect inference on TypeEvalPy and the highest TypeSim on real-world code (77% vs. 61% for GPT-4o), RightTyper produces annotations of substantially higher quality than prior approaches—enabled by grounding types in actual runtime behavior and supporting the full range of Python language features.
@@
Table 1. [RQ2] RightTyper is fast: While Type4Py and QuAC are faster, RightTyper’s execution time on the real-world code from §4.3 remains within reasonable bounds, and faster than pytype and GPT-4o, in contrast to MonkeyType, which takes nearly four hours to complete.
@@
Table 1 shows the previously achieved TypeSim scores and running times. Type4Py and QuAC are the fastest, completing in 3 and 4 minutes, respectively, followed by RightTyper at 7 minutes and pytype at 11 minutes. GPT-4o completes in 12 minutes when issuing four prompts asynchronously; increased parallelism could likely reduce its execution time. MonkeyType is by far the slowest, requiring nearly four hours. We conclude that RightTyper is fast enough for practical use.

### v3

@@ Abstract.
This paper presents RightTyper, a novel hybrid approach for Python that produces accurate and precise type annotations grounded in actual program behavior. RightTyper grounds inference in types observed during actual program execution and combines these observations with static analysis and name resolution to produce substantially higher-quality type annotations than prior approaches. Through principled, statistically guided adaptive sampling, RightTyper balances runtime overhead with the need to observe sufficient execution behavior to infer high-quality type annotations. We evaluate RightTyper against static, dynamic, and AI-based systems on both synthetic benchmarks and real-world code, and find that it consistently achieves higher semantic similarity to ground-truth and developer-written annotations, respectively, while incurring only approximately 27% runtime overhead.
@@ 1.1. Contributions
This paper proposes RightTyper, a novel hybrid type inference approach for Python. RightTyper sidesteps pitfalls of static and AI-based methods by sampling types observed during actual program execution. It employs adaptive sampling strategies to minimize runtime overhead while maintaining inference accuracy. RightTyper monitors function calls using a Poisson process, in which observations occur only during brief, randomly timed capture windows, allowing most instrumentation to be disabled outside these windows. For container types, RightTyper fully scans small containers, but applies Good–Turing estimation as a stopping criterion for larger ones, an estimator developed by Turing to estimate unseen events (good-turing; vanderWalt2021). Additionally, RightTyper tracks previously seen containers to avoid re-typing them, while continuing to sample to detect changes. Together, these strategies achieve low overhead (overall $\sim$27%) while using statistically principled sampling to obtain representative coverage.
@@
  • Introduces RightTyper, a novel dynamic–static hybrid approach to type inference for Python (§2);
  • Presents an open-source prototype implementation of RightTyper3);
  • Provides an enhanced implementation of TypeSim—a semantic type similarity metric from prior work (typybench)—addressing limitations in the original and released publicly (§4.1);
  • Demonstrates, through extensive empirical evaluation, that RightTyper produces substantially higher-quality type annotations than prior approaches while incurring significantly lower runtime overhead (§4.2, §4.3, §4.4);
  • Shows that RightTyper’s adaptive sampling strategies achieve high type recall while substantially reducing sampling effort (§4.5, §4.6).
@@ 2.1. Instrumentation and Event Sampling
Most of RightTyper’s instrumentation utilizes sys.monitoring, a low-overhead, fine-grained mechanism introduced in Python 3.12 for tracking various events during execution (pep669). This feature allows selective activation and deactivation of monitoring for specific portions of user code. When sys.monitoring does not provide the required events, RightTyper inserts additional instrumentation (§3).
@@ 2.3. Typing Generics
For large containers, after collecting a minimum number of samples (default 24), RightTyper computes the singleton ratio $\hat{r}$—the proportion of observed types that appear exactly once:
@@
where $C$ is a counter mapping each observed type $t$ to its count $C(t)$, $n$ is the number of samples, and $\mathbf{1}[\cdot]$ is the indicator function. When this ratio falls below a threshold $\tau$ (default $\tau=0.05$) for all parameters (e.g., dictionary key and value types are counted separately), sampling terminates. A hard cap (default 128 samples) further prevents pathological cases from dominating execution time. RightTyper aggregates all types observed across inspections into a cumulative history, from which it generates annotations.
@@ 2.6. Simplifying Types
Despite the simplifications above, annotations can still be complex. RightTyper offers options to trade precision for readability, such as collapsing repeated argument types (e.g., tuple[int, int, int] to tuple[int, ...]), limiting nesting depth, and capping union size (unions exceeding a configurable limit are collapsed to Any).
@@ 3. Implementation
RightTyper uses sys.monitoring events (PY_START, PY_RETURN, PY_YIELD, PY_UNWIND, and CALL) to capture function arguments, return values, yielded values, exception-driven unwinding, and call-site information. Yielding temporarily suspends execution to produce a value before later resuming. Generator functions and coroutines (including async functions) both yield values, but coroutines may also receive values during execution; because sys.monitoring does not expose these received values, RightTyper supplements monitoring with AST-based instrumentation, intercepting code loading to inject the necessary hooks before execution.
@@
Field Classes
Classes defined with @dataclass, attrs, or NamedTuple have auto-generated __init__ (or __new__) methods whose source code is unavailable at instrumentation time. RightTyper detects instantiation of such classes via the CALL event, dynamically enables monitoring on their constructors, and maps observed argument types to field annotations.
@@ 4. Evaluation
Do type annotations generated by RightTyper improve upon those produced by prior approaches? (§4.2, §4.3)
@@
How does RightTyper’s runtime overhead compare with that of other dynamic approaches? (§4.4)
RQ3::
How effective is RightTyper’s sampling approach? (§4.5)
RQ4::
How effective is Good–Turing estimation as a stopping criterion for container sampling? (§4.6)
@@ 4.1. The TypeSim Metric and Its Implementation
Because the original implementation builds on mypy, the types it compares reflect not only written annotations but also types inferred by mypy from context, conflating annotation accuracy with inference behavior. In addition, it compares only attributes declared directly in each class body rather than the full set of inherited attributes, which can yield a similarity of zero between a class and its parent. To avoid these issues, we reimplement TypeSim using libcst and operate directly on annotation syntax.
We further extend TypeSim to support commonly used typing constructs—such as Callable, TypedDict, Literal, and type aliases—that the original implementation ignored and therefore excluded from evaluation. Our version evaluates all of these constructs: callable types are compared structurally, literal types receive partial credit when they share a base type, and type aliases and type variables are expanded during extraction rather than rejected at comparison time. We also normalize semantically equivalent constructs before comparison—for example, Generator[X, None, None] is reduced to Iterator[X]—so that superficial differences in spelling do not penalize otherwise correct predictions.
As in the original TypeSim, and following prior work, we exclude top-level Any annotations from evaluation because they are uninformative (10.1145/3689783; typybench). For consistency, we also exclude top-level object annotations and assign a score of zero to Any and object when they appear within compound types (e.g., dict[str, Any]). The original implementation could assign full credit in such cases through string matching, even when part of the type structure was uninformative.
@@
Figure 10(a) presents the results in terms of TypeSim scores. With an overall 99.8% TypeSim, RightTyper achieves a near-perfect score, with only two cases failing to yield an exact match. In one case, the benchmark requests the type of an abstract method that does not execute. In the other case, RightTyper correctly types a zip iterator using the Iterator protocol—necessary to specify element types, since zip itself does not support them—but receives a lower score. This lower score occurs because TypeSim penalizes protocol-based annotations: protocols expose only the minimal set of required operations and therefore appear less similar to richer concrete types under the metric. GPT-4o, at 95%, also performs well, only incorrectly inferring “Unknown” or Any in a few cases. It is followed by pytype at 60.6%; it suffers from low coverage, consistent with prior work showing that its conservative static analysis leads to limited coverage (10.1145/3652153). MonkeyType and QuAC achieve 32.9% and 23%, respectively, in large part because they lack support for inferring variable types (Figure 11). QuAC operates by collecting sets of required attributes and using BM25—a term-frequency ranking function from information retrieval—to match these sets to candidate types (bm25). It is therefore plausible that the benchmark’s short programs—designed to facilitate manual ground-truth verification—provide limited context for this matching process, reducing QuAC’s effectiveness (10.1145/3689783). Similarly, Type4Py’s relatively low performance (24.8%) may be due to the benchmark programs likely representing out-of-distribution inputs for its model.
@@
(b) Real-world code (10,964 annotations)
Figure 10. [RQ1] RightTyper achieves the highest type inference accuracy: (a) On TypeEvalPy, RightTyper achieves 99.8% TypeSim, outperforming GPT-4o (95%). (b) On real-world code, RightTyper achieves 77.2% TypeSim, outperforming GPT-4o (72.6%) and MonkeyType (57.5%).
@@
Although the programs in TypeEvalPy provide broad coverage of Python language features, they are significantly simpler than, and differ markedly from, commonly deployed Python code. We therefore next evaluate RightTyper on the real-world Python packages black, flask, httpx, pylint, and rich, which together contain over 10,000 annotations. We select these packages based on several criteria. First, all five are well-annotated open-source Python projects, with annotations developed and refined by expert maintainers over many years of use with type checkers—making them ideal ground-truth sources. Second, they are complex, production-quality codebases that exercise diverse Python features including inheritance hierarchies, generic types, protocols, decorators, context managers, and AST manipulation. Third, they are widely used (pepy.tech reports billions of downloads) and well-maintained, reducing the risk of ground-truth errors. Finally, the packages span command-line tools, libraries, and a web framework, allowing us to drive them end to end with varied inputs, achieving broader coverage than test suites alone. As ground truth, we use the projects’ original type annotations; that is, we evaluate the semantic similarity of inferred annotations to developer-written ones using TypeSim.
@@
Prior to evaluating each system, we remove all type annotations from the source code, except for special constructs such as dataclasses and NamedTuple where removing annotations would alter program semantics; in these cases, we replace the annotations with a type alias placeholder, allowing us to distinguish them from tool-inferred types during evaluation. We exclude annotations originally typed as Any, since they carry no type information and cannot meaningfully distinguish tool quality. We also remove any docstrings, as these can influence AI-based type inference; this ensures a level comparison with non-AI methods that cannot leverage natural language hints.
@@
When evaluating RightTyper and MonkeyType, we execute each target package multiple times with varying options and inputs—keeping them identical across the two systems—and also run their test suites to further exercise production code, thereby collecting typing information across a range of execution scenarios. RightTyper excludes test functions from tracing by default; we configure MonkeyType to do the same, so that only production code is typed. As before (§4.2), we run MonkeyType without sampling, allowing it to record every execution and achieve its best possible results. Rather than having MonkeyType annotate the source code directly, we configure it to generate stub files, as in-source annotation causes MonkeyType to emit import statements that introduce circular dependencies. Applying MonkeyType and QuAC to these sources uncovered several bugs, including an assertion failure in QuAC and syntax errors in MonkeyType’s output; we corrected these issues. To include LLM-based results, we reimplement querying GPT-4o using TypyBench’s “single-file context” prompt, as its replication package does not provide an implementation.
@@
Figure 10(b) presents the results. With an overall TypeSim score of 77.2%, RightTyper achieves the highest similarity to developer-annotated types. It is followed by GPT-4o and MonkeyType, which at 72.6% and 57.5% also show good performance, and at a considerable distance by Type4Py at 29.6%, QuAC at 28%, and pytype at 19.8%.
@@
Figure 11. [RQ1] Detailed results: (a) TypeSim scores by annotation category show RightTyper excels on both functions and variables, while QuAC and MonkeyType cannot infer variable types. (b) On real-world code, RightTyper achieves the highest TypeSim (77.2%) and Coverage (93.1%).
@@
RightTyper exhibits some similar limitations, including occasional overly concrete typing and missed coverage, the latter being expected given that it executes the same scenarios as MonkeyType. However, RightTyper achieves higher scores overall due to more complete language support—such as inferring variable types, handling keyword and variadic arguments, inferring Callable argument types, and handling nested functions—all of which MonkeyType lacks—as well as its ability to better simplify inferred types, yielding cleaner annotations that more closely match the originals (Figure 11). The lower scores on variable annotations stem from variables captured only at function exit (missing intermediate types), unexecuted branches, and developers annotating with broader types than observed at runtime. Finally, in some cases RightTyper correctly emits annotations with type parameters (cf. §2.4), but TypeSim yields a zero score because it does not capture these forms of semantic equivalence.
Given RightTyper’s strong results on both TypeEvalPy and real-world code, we conclude that it improves annotation quality relative to prior approaches.
[RQ1] Summary: With near-perfect inference on TypeEvalPy and the highest TypeSim on real-world code (77.2% vs. 72.6% for GPT-4o and 57.5% for MonkeyType), RightTyper produces annotations of higher quality than prior approaches—enabled by grounding types in actual runtime behavior and supporting a wider range of Python language features.
@@
Table 1. [RQ2] RightTyper is fast: While Type4Py and QuAC are faster, RightTyper’s execution time on the real-world code from §4.3 remains within reasonable bounds, matching pytype and faster than GPT-4o, in contrast to MonkeyType, which takes over 4.5 hours to complete.
@@
Table 1 reports the previously achieved TypeSim scores alongside running times. Type4Py and QuAC are the fastest, completing in about 5 and 6 minutes respectively, followed by RightTyper at 8 minutes. pytype and GPT-4o complete in 14 and 22 minutes, respectively; for GPT-4o, increased parallelism could likely reduce its execution time. MonkeyType is by far the slowest, requiring nearly five hours. We conclude that RightTyper is fast enough for practical use.
@@
4.5. [RQ3] Sampling Effectiveness
Table 2. [RQ3] RightTyper’s trace sampling is effective: MonkeyType’s Bernoulli sampling degrades accuracy with modest speed gains, while RightTyper’s Poisson-timed sampling achieves both the highest accuracy and lowest execution time.
Metric MonkeyType RightTyper
unsampled p=0.01p{=}0.01 p=0.001p{=}0.001
TypeSim 57.5% 37.4% 22.6% 77.2%
Time (h:mm) 4:44 3:05 3:04 0:08
Total traces 126,377,033 1,463,432 151,559 218,797
Unique functions 3,869 2,318 1,391 3,895
Larger TypeSim and more functions are better; shorter times and fewer traces are better.
Event Sampling
A key design choice in RightTyper is its use of Poisson-timed event sampling to reduce overhead while preserving annotation quality (§2.1). To assess the importance of this design, we compare it against MonkeyType’s Bernoulli sampling, which records all types from a random subset of calls. We evaluate MonkeyType with sampling rates p=0.01p{=}0.01 and p=0.001p{=}0.001 on the same real-world code described in §4.3.
Table 2 reports the results, including the number of total traces (individual call observations recorded) and unique functions (distinct functions for which at least one trace was recorded). MonkeyType’s TypeSim degrades substantially under sampling, dropping from 57.5% without sampling to 37.4% at p=0.01p{=}0.01 and 22.6% at p=0.001p{=}0.001. Despite this degradation, execution time remains high—at roughly three hours for both sampling rates—suggesting that instrumentation and related overhead dominate MonkeyType’s running time in these regimes.
[…]
Container Sampling
RightTyper tracks containers and either scans or samples their contents to infer full container types (§2.3). To assess the effectiveness of this approach, we further instrument RightTyper to log every container observation—including the action taken (full scan, sampling, or spot-check), the number of samples drawn, and the types found—and to perform an exhaustive ground-truth scan of every container encountered. We then re-execute RightTyper on the same real-world code from Section 4.3 and compare the types that it finds against the ground truth.
RightTyper records 6.6M container observations, of which 97.4% involve small containers ($\leq 32$ entries) that are fully scanned. The remaining 2.6% are large containers that trigger adaptive sampling. Overall, RightTyper achieves perfect recall in 99.6% of observations and a mean recall of 99.83%. On large containers, where sampling replaces exhaustive scanning, it attains 85.2% perfect recall with a mean recall of 93.8%.
Sampling also yields substantial efficiency gains. On average, RightTyper draws 26 samples from large containers—mean size 837 entries—reducing total element inspections by 93.3% relative to exhaustive scanning—such as that performed by MonkeyType. Following the initial observation, RightTyper’s change-detection mechanism—which triggers re-sampling when a container’s size changes or when a spot check encounters a new type—detects 74.3% of ground-truth type changes with a median latency of one observation.
[…]
4.6. [RQ4] Good–Turing Estimation
Figure 13. [RQ4] Good–Turing adapts sampling effort to container diversity: it matches random k=48k{=}48 in recall while drawing 47% fewer total samples, saving on single-type containers and investing more on diverse ones. Percentages indicate container observations in that bucket. Annotations above Good–Turing bars show the mean number of samples drawn ($\bar{k}$).
To evaluate the Good–Turing stopping criterion used in container sampling (§2.3), we compare it against a baseline that draws a fixed number $k$ of random samples (with replacement) from each large container. The additional instrumentation described in Section 4.5 provides ground-truth type counts for all containers, enabling us to compute expected recall analytically. The probability of observing type $t$ at least once in $k$ draws with replacement, and the expected recall (fraction of distinct types observed), are given by
[…]
where $f_{t}$ is the frequency of type $t$, NN is the container size, and $T$ denotes the set of distinct types.
Figure 13 shows the results grouped by the number of ground-truth types per container. Good–Turing’s adaptive strategy concentrates effort where it matters most. On homogeneous containers (1 type; 68.9% of cases), all strategies achieve perfect recall; Good–Turing draws the configured minimum of 24 samples. On moderately polymorphic containers (2–10 types; 27.4%), Good–Turing performs comparably to fixed k=32k{=}32 (63.3% vs. 63.8% in aggregate recall). On highly polymorphic containers (11+ types; 3.7%), Good–Turing draws a mean of 62 samples and achieves 73.5% recall, comparable to k=64k{=}64 (73.2%). The lower mean sample count in this bucket reflects the union-size cap (§2): once the number of observed types exceeds the maximum union size, further sampling cannot improve the annotation (which collapses to Any), so RightTyper stops early. In aggregate, Good–Turing matches k=48k{=}48 in recall (89% vs. 89.5%) while drawing 47% fewer total samples (1.39 vs. 2.62 million), by avoiding over-sampling the dominant homogeneous containers.
[RQ4] Summary: Good–Turing estimation is highly effective as a stopping criterion for container sampling. In aggregate, it matches the recall of fixed k=48k{=}48 sampling while drawing 47% fewer total samples by adaptively allocating effort based on container diversity.